A publisher has a fixed cost of $ 250,000 associated with the production of a college mathematics book. The contribution to profit and fixed cost from the sale of each book is $6.25. Determine the number of books which must be sold in order to break even.
Found 2 solutions by Theo, ikleyn:
Answer by Theo(13342)

the fixed cost is 250,000
the contribution to profit and fixed cost from the sale of each book is 6.25.
let x = the number of books required to cover the fixed cost.
your equation is 250,000 = 6.25 * x
solve for x to get x = 250,000 / 6.25 = 40,000
if the publisher sells 40,000 books, he breaks even.
any number of books sold greater than 40,000 is profit as far as i know.

note that the contribution to profit is the sale price per unit minus the variable cost per unit.

for example:

if the sale price per book is 5.00 and the variable cost per book is 2.00 , then the contribution to profit is 3.00 per book.

here's a reference.

https://www.investopedia.com/terms/c/contributionmargin.asp

here's another.

https://corporatefinanceinstitute.com/resources/knowledge/accounting/contribution-margin-overview/








Answer by ikleyn(52879)

.
A publisher has a fixed cost of $ 250,000 associated with the production of a college mathematics book.
The contribution to profit and fixed cost from the sale of each book is $6.25.
Determine the number of books which must be sold in order to break even.
~~~~~~~~~~~~~~~~


Formulation of the problem in the post is TOTALLY INCORRECT.

In order for the problem makes sense, the selling price for each book should be given.


In the post, the problem is presented absolutely UNPROFESSIONALLY with using wrong terminology.